OpenStudy (anonymous):

Given f(x) and g(x), approximate the following.(Round your answer to three decimal places.) f(x) = x^3/(x^2+x+3) g(x) = ((sqrt5x) -x^3)^(3/2) Find ((f + g)(1.14) - (f/g)(1.14)) / ([ (f o f)(5.6) ]^2) (I tried this multiple times but I can't seem to get the right answer)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

(f o f)(5.6) ]^2)

OpenStudy (anonymous):

^ sure thats correct

OpenStudy (anonymous):

its just calculator work, make sure you use brackets

OpenStudy (anonymous):

I cant exactly help you with calculator work over the net

OpenStudy (anonymous):

(f+g) = f(x) +g(x) (f/g)(x) = f(x) / g(x) , when g(x) not zero etc

OpenStudy (anonymous):

its fairly easy things to remember

OpenStudy (anonymous):

The denominator is ((f of f) (5.6))^2

OpenStudy (anonymous):

ok so that means its f ( f(5.6) )

OpenStudy (anonymous):

I know how to do it but for some reason the answer is never right

OpenStudy (anonymous):

so go to the expression for f , sub in 5.6 , then whatever you answer is , sub it back into f , and then sqaure whatever you get

OpenStudy (anonymous):

yeah and then then the completed (f of f) answer is squared

OpenStudy (anonymous):

its calculator work , just use brackets for everything

OpenStudy (anonymous):

kk